# Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands

### v1

@@ IV S4: Short-range Nature of the $\Delta\sim\Delta\ln(W/2\Delta)$ Crossover
In Fig. 7(b), the $D_{s}$ curves for both geometric and conventional contribution reach maximum values as $\Delta$ approaches some magnitude between $E_{g}$ and $E_{g}+W$. This phenomenon is due to the inter-band reduction effect [46], because the usual $D_{s}\sim\Delta$ linear relation for flat bands cannot tend to infinity in a multi-band scenario; as $\Delta$ approaches the energy of other high-energy bands, the geometry of bands within the interaction scale gets trivialized and $D_{s}$ starts to decrease after the $2\Delta$ interaction window finishes including the high-energy bands. This geometry trivialization also explains in Fig. 4 why $D_{s}^{\text{geo}}$ decreases slightly when $E_{g}$ keeps closing after reaching $\Delta$.
V S5: The Uniform Pairing Condition at Zero-temperature for Chiral Models
@@
Take the bipartite Lieb lattice model Eq. 29 as an example. There are two limits: (1) the gap-closed limit $t\delta\ll\Delta$ and (2) the isolated limit $t\delta\gg\Delta$. For the former, although the flat band has no component from orbital $B$ (see Fig. 6(b)), the two dispersive bands that have entered the interaction scale do contain orbital $B$. Therefore, if one wants to impose UPC and correctly consider the inter-band effect between the two dispersive bands, $U_{B}/U_{A}$ must be tuned as $\delta$ changes. When $\delta$ increases from 0 to $\delta\sim\Delta/t$, the $B$-orbital-resolved density of states in the interaction window vanishes, implying $U_{B}\rightarrow\infty$. This shows that in the regime $0\leq\delta\lesssim\Delta/t$, UPC is not a good approximation if one fixes $U_{B}$. However, once the isolated limit is reached, orbital $B$ has been removed from the interaction scale, then the value of $U_{B}$ will not affect the superfluid weight. Therefore, in the regime $t\delta\gg\Delta$, UPC is always well established as long as $U_{A}=U_{C}$, as pointed out by Ref. [16].
@@
Here we solve the self-consistency gap equations for the chiral models Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands and 3, to point out in what parameter regime the UPC is a good approximation if $U_{\alpha}$ are fixed. Unlike the bipartite Lieb lattice, in the three-band chiral model, as the band gap is opened by the mass $m$, some orbital $B$ component shifts from the dispersive bands to the flat band. Therefore, it is never removed from the interaction scale.
@@
Figure 8: $U_{\alpha}$ as function of $E_{g}$ solved from UPC $\Delta_{\alpha}\equiv\Delta$ for the chiral models with $J=\zeta_{J}=k_{\Lambda}=1$. Blue curves: $\Delta=0.01\varepsilon_{\Lambda}$; Red curves: $\Delta=0.1\varepsilon_{\Lambda}$. (a) for two-band model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands, $E_{g}=2m$, $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}$; (b)(c) for case (i),(ii) of the three-band model Eq. 3, respectively, with $E_{g}=m$, $\varepsilon_{\Lambda}=\zeta_{J}k_{\Lambda}^{J}$.
@@
In Fig. 8, we impose UPC by fixing all $\Delta_{\alpha}\equiv\Delta$ and solve $U_{\alpha}$ as a function of $E_{g}$ from the gap equation:
$1=\frac{U_{\alpha}}{N}\sum_{\textbf{k},l}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}},$ (34)
@@
Where k is summed over the cutoff region $k<k_{f}=1$, $l$ is summed over all the two or three bands of the model, and $\mu$ is set to be 0. For the two-band and three-band chiral model, the variation of the ratio $U_{A}/U_{B}$ with $E_{g}$ tells whether UPC is a good approximation in some regime (for the three-band model we set $U_{A}=U_{C}$). In Fig. 8(a), we have $U_{A}=U_{B}$ at $E_{g}=0$ because the two orbitals are symmetric; as the gap opens, $U_{A},U_{B}$ increase because the density of both orbitals near $\mu=0$ are lowered. Compared to the $\Delta=0.1\varepsilon_{\Lambda}$ case (red curves), we find the change of $U_{A}/U_{B}$ with $E_{g}$ is smaller for $\Delta=0.01\varepsilon_{\Lambda}$ (blue curves). Therefore, for the same regime $50<E_{g}/\Delta<$ 5, UPC is a better approximation for smaller $\Delta$. Fig. 8(b)(c) can be similarly analyzed. We find for case (i) of the three-band model ($C(m,k)=1$, Fig. 8(b)), UPC is again better for the smaller $\Delta$, whereas for case (ii) of the three-band model with $c(m)=m$ (Fig. 8(c)), UPC is better for larger $\Delta$ (red curves) in the entire regime $50<E_{g}/\Delta<$ 5.
@@ VI S6: Stability of the Uniform Pairing Condition with Temperature
We analytically show the uniform pairing condition (UPC) for flat bands in proximity to dispersive bands is stable against the change of temperature if the interaction is weak (i.e., either $\Delta<E_{g}$ or $\Delta>E_{g}$ but $\Delta-E_{g}\ll W_{d}$, both of which lead to much smaller electron density in dispersive bands than flat bands).
@@
$\Delta_{\alpha}(T)=-\frac{U_{\alpha}}{N}\sum_{\textbf{k},ll^{\prime}}u_{l\textbf{k}\alpha}^{*}u_{l^{\prime}\textbf{k}\alpha}\langle c_{l,-\textbf{k}\downarrow}c_{l^{\prime}\textbf{k}\uparrow}\rangle,$ (35)
@@
Let’s choose parameters $U_{\alpha}$ such that UPC is satisfied at $T=0$; we further make the assumption that it also holds for all finite $T$ below the transition temperature, i.e., $\Delta_{1}(T)=...=\Delta_{s}(T)=\Delta(T)$ (this assumption will be checked in the following calculations). With this assumption, Eq. 35 simplifies to
$1=\frac{U_{\alpha}}{N}\sum_{\textbf{k},l}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2},$ (36)
where $E_{l\textbf{k}}=\sqrt{\xi_{l\textbf{k}}^{2}+\Delta(T)^{2}}$. Now that the $s$ equations of Eq. 36 are decoupled, solving $\Delta(T)$ from them individually gives a curve $\widetilde{\Delta}_{\alpha}(T)$. Then, our assumption can be proved if these $s$ curves coincide.
Choosing any $\alpha$th equation from Eq. 36, it reads
$1=\frac{U_{\alpha}}{N}\sum_{\textbf{k}}\bigg{(}\sum_{l\in\text{flat}}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2}+\sum_{l^{\prime}\in\text{disp}}\frac{|u_{l^{\prime}\textbf{k}\alpha}|^{2}}{2E_{l^{\prime}\textbf{k}}}\tanh\frac{\beta E_{l^{\prime}\textbf{k}}}{2}\bigg{)},$ (37)
where $l$ is summed over all the degenerate flat bands and $l^{\prime}$ is over all dispersive bands. Under the condition $\mu\approx 0$ and $\Delta(0)<E_{g}$ or $\Delta(0)>E_{g}$ but $\Delta(0)-E_{g}\ll W_{d}$, the r.h.s. of Eq. 37 is mainly contributed by the first term because these from the dispersive bands are cut off by factor $1/E_{l^{\prime}\textbf{k}}$ in most regions of k space. This means Eq. 37 can be approximated as
$1\approx\frac{U_{\alpha}}{N}\sum_{\textbf{k},l\in\text{flat}}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2},\,\,\,\forall T.$ (38)
Combining the Eq. 38 at $T=0$ and $T=T_{MF}$ (defined from $\widetilde{\Delta}_{\alpha}(T_{MF})=0$), it leads to
$\sum_{\textbf{k},l\in\text{flat}}|u_{l\textbf{k}\alpha}|^{2}\bigg{[}\frac{1}{2T_{MF}}-\frac{1}{\Delta(0)}\bigg{]}\approx 0,$ (39)
@@
In Fig. 9, we calculated the $\widetilde{\Delta}_{\alpha}(T)$ curves for the two-band chiral model. For this model, if the band gap completely closes ($m=0$), then the two orbitals are symmetric, and $\Delta_{1}(T)=\Delta_{2}(T)$ holds for any temperature once we set $U_{1}=U_{2}$. Therefore, we are interested in the case $m\neq 0$. As long as the mass $m$ is not too large, both orbitals $\alpha=1,2$ contribute to the flat band, so the curves show the UPC holds well until $T$ gets very close to $T_{MF}$.
Figure 9: Plot of $\widetilde{\Delta}_{\alpha}(T)$ ($\alpha=1,2$) solved from Eq. 36 for the two-band chiral model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands with $J=1$, $\zeta_{J}=1$. $m$ is fixed to be $0.2W_{d}$ so $E_{g}=2m=0.4W_{d}$. Blue curves for the case $\Delta(0)<E_{g}$; orange curves for the case $\Delta(0)>E_{g}$ but $\Delta(0)-E_{g}\ll W_{d}$.
@@ VII S7: Temperature Dependence of Superfluid Weight and BKT Transition
$\Delta_{\alpha}=\frac{U_{\alpha}}{N}\sum_{\textbf{k},l}|u_{l\textbf{k}\alpha}|^{2}\frac{\bar{\Delta}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2}+O(\delta\Delta_{\alpha}),$ (40)
@@
where $\bar{\Delta}=(1/s)\sum_{\alpha=1}^{s}\Delta_{\alpha}$ ($s=2$ or 3) is the averaged order parameter, $\delta\Delta_{\alpha}=\Delta_{\alpha}-\bar{\Delta}$ and $E_{l\textbf{k}}=\sqrt{\xi_{l\textbf{k}}^{2}+\bar{\Delta}^{2}}$. Eq. 40 is a good approximation of the exact equation 35 when the pairing matrix $\hat{\Delta}=\text{diag}\{\Delta_{1},...,\Delta_{s}\}$ is close to uniform, with the inter-band $O(\delta\Delta_{\alpha})$ terms neglected. Its solution tells the lowest order how $\hat{\Delta}$ deviates from uniform pairing.
@@
Figure 10: $\bar{\Delta}$ and $\Delta_{A}/\Delta_{B}$ solved for the flattened three-band model, with $c(m)=m$. (a)(b) Fixing $U_{A}=U_{B}=2.51\varepsilon_{\Lambda}$ such that UPC is satisfied at $E_{g}=T=0$, with $\Delta_{\alpha}(0,0)=0.1\varepsilon_{\Lambda}$; (c)(d) similarly, fixing $U_{A}=U_{B}=12.56\varepsilon_{\Lambda}$ such that $\Delta_{\alpha}(0,0)=0.5\varepsilon_{\Lambda}$.
As an example, in Fig. 10, we plot the $\Delta_{\alpha}(E_{g},T)$ solved from Eq. 40 for the flattened three-band chiral model (case (ii)) by taking $c(m)=m$. In the calculation, $U_{A},U_{B}$ are fixed such that the UPC is satisfied at $E_{g}=0,T=0$. In Fig. 10(a)(b), the UPC $\Delta_{\alpha}(0,0)=0.1\varepsilon_{\Lambda}$, we find in (a) as $E_{g}=m$ closes, $\bar{\Delta}$ increases due to the enhanced density of states, so $T_{MF}$ (the temperature at which $\bar{\Delta}$ decrease to 0) also increases. (b) shows the ratio of $\Delta_{A}/\Delta_{B}$ which tells how uniform the pairing is. At $T=0$, as $m$ increases, the pairing becomes nonuniform at $E_{g}/\varepsilon_{\Lambda}=0.2$ (i.e., $E_{g}/\bar{\Delta}\approx 2$) and becomes uniform again at $E_{g}/\varepsilon_{\Lambda}=0.4$ ($E_{g}/\bar{\Delta}\approx 4$), which reproduces the red curves in Fig. 8(c). This non-uniformity with $E_{g}$ becomes less severe as $T$ increases. In Fig. 10(c)(d) we show the solutions with UPC $\Delta_{\alpha}(0,0)=0.5\varepsilon_{\Lambda}$ (which was not shown in Fig. 8(c)). In (d), we find the UPC holds well except for the regime $E_{g}/\varepsilon_{\Lambda}>1.5$, $T/\varepsilon_{\Lambda}<0.2$, as the model approaches the atomic limit.
After obtaining $\Delta_{\alpha}(T)$, we use UPC as an approximation and insert $\bar{\Delta}(T)$ into Eq. II to compute $D_{s}(T)$. From this, one can further calculate the BKT transition temperature $T_{c}$, as shown in Fig. 4(b) and Fig. 5(b). Fig. 11 shows the temperature-dependence of $D_{s}$ using the example of the two-band chiral model. The $D_{s}(T)$ curves are very flat at $T<0.3T_{MF}$ and have a finite slope near $T_{MF}$, which is a general feature of flat bands with $s$-wave uniform order parameter. $T_{c}$ is extracted from the $D_{s}(T)$ curves by their intersections with the universal line of $(8/\pi)T$. When $D_{s}(0)/\Delta(0)$ is small, the intersections lie in the flat part of the curves and give $T_{c}\approx(\pi/8)D_{s}(0)$, exhibiting a Uemura relation [51].
@@
Figure 11: $D_{s}$ vs temperature for the two-band model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands. Different colors show $E_{g}=0.5\varepsilon_{\Lambda}$ (green), $E_{g}=\Delta(0)$ (orange) and $E_{g}=0$ (blue). (a) $\Delta(0)=0.1\varepsilon_{\Lambda}$, (b) $\Delta(0)=0.01\varepsilon_{\Lambda}$.

### v2

@@ IV S4: Short-range Nature of the $\Delta\sim\Delta\ln(W/2\Delta)$ cross-over
In Fig. S2(b), the $D_{s}$ curves for both geometric and conventional contribution reach maximum values as $\Delta$ approaches some magnitude between $E_{g}$ and $E_{g}+W$. This phenomenon is due to the inter-band reduction effect [34], because the usual $D_{s}\sim\Delta$ linear relation for flat bands cannot tend to infinity in a multi-band scenario; as $\Delta$ approaches the energy of other high-energy bands, the geometry of bands within the interaction scale gets trivialized and $D_{s}$ starts to decrease after the $2\Delta$ interaction window finishes including the high-energy bands. This geometry trivialization also explains in Fig. 4 of MT why $D_{s}^{\text{geo}}$ decreases slightly when $E_{g}$ keeps closing after reaching $\Delta$.
V S5: Uniform Pairing Condition at Zero-temperature for Chiral Models
@@
Take the bipartite Lieb lattice model Eq. (S22) as an example. There are two limits: (1) the gap-closed limit $t\delta\ll\Delta$ and (2) the isolated limit $t\delta\gg\Delta$. For the former, although the flat band has no component from orbital $B$ (see Fig. S1(b)), the two dispersive bands that have entered the interaction scale do contain orbital $B$. Therefore, if one wants to impose UPC and correctly consider the inter-band effect between the two dispersive bands, $U_{B}/U_{A}$ must be tuned as $\delta$ changes. When $\delta$ increases from 0 to $\delta\sim\Delta/t$, the $B$-orbital-resolved density of states in the interaction window vanishes, implying $U_{B}\rightarrow\infty$. This shows that in the regime $0\leq\delta\lesssim\Delta/t$, UPC is not a good approximation if one fixes $U_{B}$. However, once the isolated limit is reached, orbital $B$ has been removed from the interaction scale, then the value of $U_{B}$ will not affect the superfluid weight. Therefore, in the regime $t\delta\gg\Delta$, UPC is always well established as long as $U_{A}=U_{C}$, as pointed out by Ref. [20].
@@
Here we solve the self-consistency gap equations for the chiral models Eq. 1 and 3 of MT, to point out in what parameter regime the UPC is a good approximation if $U_{\alpha}$ are fixed. Unlike the bipartite Lieb lattice, in the three-band chiral model, as the band gap is opened by the mass $m$, some orbital $B$ component shifts from the dispersive bands to the flat band. Therefore, it is never removed from the interaction scale.
@@
Figure S3: $U_{\alpha}$ as function of $E_{g}$ solved from UPC $\Delta_{\alpha}\equiv\Delta$ for the chiral models with $J=\zeta_{J}=k_{\Lambda}=1$. Blue curves: $\Delta=0.01\varepsilon_{\Lambda}$; Red curves: $\Delta=0.1\varepsilon_{\Lambda}$. (a) for two-band model Eq. 1 of MT, $E_{g}=2m$, $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}$; (b)(c) for case (i),(ii) of the three-band model Eq. 3 of MT, respectively, with $E_{g}=m$, $\varepsilon_{\Lambda}=\zeta_{J}k_{\Lambda}^{J}$.
@@
In Fig. S3, we impose UPC by fixing all $\Delta_{\alpha}\equiv\Delta$ and solve $U_{\alpha}$ as a function of $E_{g}$ from the gap equation:
$1=\frac{U_{\alpha}}{N}\sum_{\textbf{k},l}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}},$ (S27)
@@
Where k is summed over the cutoff region $k<k_{f}=1$, $l$ is summed over all the two or three bands of the model, and $\mu$ is set to be 0. For the two-band and three-band chiral model, the variation of the ratio $U_{A}/U_{B}$ with $E_{g}$ tells whether UPC is a good approximation in some regime (for the three-band model we set $U_{A}=U_{C}$). In Fig. S3(a), we have $U_{A}=U_{B}$ at $E_{g}=0$ because the two orbitals are symmetric; as the gap opens, $U_{A},U_{B}$ increase because the density of both orbitals near $\mu=0$ are lowered. Compared to the $\Delta=0.1\varepsilon_{\Lambda}$ case (red curves), we find the change of $U_{A}/U_{B}$ with $E_{g}$ is smaller for $\Delta=0.01\varepsilon_{\Lambda}$ (blue curves). Therefore, for the same regime $50<E_{g}/\Delta<$ 5, UPC is a better approximation for smaller $\Delta$. Fig. S3(b)(c) can be similarly analyzed. We find for case (i) of the three-band model ($C(m,k)=1$, Fig. S3(b)), UPC is again better for the smaller $\Delta$, whereas for case (ii) of the three-band model with $c(m)=m$ (Fig. S3(c)), UPC is better for larger $\Delta$ (red curves) in the entire regime $50<E_{g}/\Delta<$ 5.
@@ VI S6: Stability of the Uniform Pairing Condition with Temperature
We analytically show the uniform pairing condition (UPC) for flat bands in proximity to dispersive bands is stable against the change of temperature if the interaction is weak (i.e., either $\Delta<E_{g}$ or $\Delta>E_{g}$ but $\Delta-E_{g}\ll W_{d}$, both of which lead to much smaller electron density in the dispersive bands than the flat bands).
@@
$\Delta_{\alpha}(T)=-\frac{U_{\alpha}}{N}\sum_{\textbf{k},ll^{\prime}}u_{l\textbf{k}\alpha}^{*}u_{l^{\prime}\textbf{k}\alpha}\langle c_{l,-\textbf{k}\downarrow}c_{l^{\prime}\textbf{k}\uparrow}\rangle,$ (S28)
@@
Let’s choose parameters $U_{\alpha}$ such that UPC is satisfied at $T=0$; we further make the assumption that it also holds for all finite $T$ below the transition temperature, i.e., $\Delta_{1}(T)=...=\Delta_{s}(T)=\Delta(T)$ (this assumption will be checked in the following calculations). With this assumption, Eq. (S28) simplifies to
$1=\frac{U_{\alpha}}{N}\sum_{\textbf{k},l}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2},$ (S29)
where $E_{l\textbf{k}}=\sqrt{\xi_{l\textbf{k}}^{2}+\Delta(T)^{2}}$. Now that the $s$ equations of Eq. (S29) are decoupled, solving $\Delta(T)$ from them individually gives a curve $\widetilde{\Delta}_{\alpha}(T)$. Then, our assumption can be proved if these $s$ curves coincide.
Figure S4: Plot of $\widetilde{\Delta}_{\alpha}(T)$ ($\alpha=1,2$) solved from Eq. (S29) for the two-band chiral model Eq. 1 of MT with $J=1$, $\zeta_{J}=1$. $m$ is fixed to be $0.2W_{d}$ so $E_{g}=2m=0.4W_{d}$. Blue curves for the case $\Delta(0)<E_{g}$; orange curves for the case $\Delta(0)>E_{g}$ but $\Delta(0)-E_{g}\ll W_{d}$.
Choosing any $\alpha$th equation from Eq. (S29), it reads
$1=\frac{U_{\alpha}}{N}\sum_{\textbf{k}}\bigg{(}\sum_{l\in\text{flat}}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2}+\sum_{l^{\prime}\in\text{disp}}\frac{|u_{l^{\prime}\textbf{k}\alpha}|^{2}}{2E_{l^{\prime}\textbf{k}}}\tanh\frac{\beta E_{l^{\prime}\textbf{k}}}{2}\bigg{)},$ (S30)
where $l$ is summed over all the degenerate flat bands and $l^{\prime}$ is over all dispersive bands. Under the condition $\mu\approx 0$ and $\Delta(0)<E_{g}$ or $\Delta(0)>E_{g}$ but $\Delta(0)-E_{g}\ll W_{d}$, the r.h.s. of Eq. (S30) is mainly contributed by the first term because these from the dispersive bands are cut off by factor $1/E_{l^{\prime}\textbf{k}}$ in most regions of k space. This means Eq. (S30) can be approximated as
$1\approx\frac{U_{\alpha}}{N}\sum_{\textbf{k},l\in\text{flat}}\frac{|u_{l\textbf{k}\alpha}|^{2}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2},\,\,\,\forall T.$ (S31)
Combining the Eq. (S31) at $T=0$ and $T=T_{MF}$ (defined from $\widetilde{\Delta}_{\alpha}(T_{MF})=0$), it leads to
$\sum_{\textbf{k},l\in\text{flat}}|u_{l\textbf{k}\alpha}|^{2}\bigg{[}\frac{1}{2T_{MF}}-\frac{1}{\Delta(0)}\bigg{]}\approx 0,$ (S32)
@@
In Fig. S4, we calculated the $\widetilde{\Delta}_{\alpha}(T)$ curves for the two-band chiral model. For this model, if the band gap completely closes ($m=0$), then the two orbitals are symmetric, and $\Delta_{1}(T)=\Delta_{2}(T)$ holds for any temperature once we set $U_{1}=U_{2}$. Therefore, we are interested in the case $m\neq 0$. As long as the mass $m$ is not too large, both orbitals $\alpha=1,2$ contribute to the flat band, so the curves show the UPC holds well until $T$ gets very close to $T_{MF}$.
@@ VII S7: Temperature Dependence of Superfluid Weight and BKT Transition
$\Delta_{\alpha}=\frac{U_{\alpha}}{N}\sum_{\textbf{k},l}|u_{l\textbf{k}\alpha}|^{2}\frac{\bar{\Delta}}{2E_{l\textbf{k}}}\tanh\frac{\beta E_{l\textbf{k}}}{2}+O(\delta\Delta_{\alpha}),$ (S33)
@@
where $\bar{\Delta}=(1/s)\sum_{\alpha=1}^{s}\Delta_{\alpha}$ ($s=2$ or 3) is the averaged order parameter, $\delta\Delta_{\alpha}=\Delta_{\alpha}-\bar{\Delta}$ and $E_{l\textbf{k}}=\sqrt{\xi_{l\textbf{k}}^{2}+\bar{\Delta}^{2}}$. Eq. (S33) is a good approximation of the exact equation (S28) when the pairing matrix $\hat{\Delta}=\text{diag}\{\Delta_{1},...,\Delta_{s}\}$ is close to uniform, with the inter-band $O(\delta\Delta_{\alpha})$ terms neglected. Its solution tells the lowest order how $\hat{\Delta}$ deviates from uniform pairing.
@@
Figure S5: $\bar{\Delta}$ and $\Delta_{A}/\Delta_{B}$ solved for the flattened three-band model, with $c(m)=m$. (a)(b) Fixing $U_{A}=U_{B}=2.51\varepsilon_{\Lambda}$ such that UPC is satisfied at $E_{g}=T=0$, with $\Delta_{\alpha}(0,0)=0.1\varepsilon_{\Lambda}$; (c)(d) similarly, fixing $U_{A}=U_{B}=12.56\varepsilon_{\Lambda}$ such that $\Delta_{\alpha}(0,0)=0.5\varepsilon_{\Lambda}$.
@@
Figure S6: $D_{s}$ vs temperature for the two-band model Eq. 1 of MT. Different colors show $E_{g}=0.5\varepsilon_{\Lambda}$ (green), $E_{g}=\Delta(0)$ (orange) and $E_{g}=0$ (blue). (a) $\Delta(0)=0.1\varepsilon_{\Lambda}$, (b) $\Delta(0)=0.01\varepsilon_{\Lambda}$.
As an example, in Fig. S5, we plot the $\Delta_{\alpha}(E_{g},T)$ solved from Eq. (S33) for the flattened three-band chiral model (case (ii)) by taking $c(m)=m$. In the calculation, $U_{A},U_{B}$ are fixed such that the UPC is satisfied at $E_{g}=0,T=0$. In Fig. S5(a)(b), the UPC $\Delta_{\alpha}(0,0)=0.1\varepsilon_{\Lambda}$, we find in (a) as $E_{g}=m$ closes, $\bar{\Delta}$ increases due to the enhanced density of states, so $T_{MF}$ (the temperature at which $\bar{\Delta}$ decrease to 0) also increases. (b) shows the ratio of $\Delta_{A}/\Delta_{B}$ which tells how uniform the pairing is. At $T=0$, as $m$ increases, the pairing becomes nonuniform at $E_{g}/\varepsilon_{\Lambda}=0.2$ (i.e., $E_{g}/\bar{\Delta}\approx 2$) and becomes uniform again at $E_{g}/\varepsilon_{\Lambda}=0.4$ ($E_{g}/\bar{\Delta}\approx 4$), which reproduces the red curves in Fig. S3(c). This non-uniformity with $E_{g}$ becomes less severe as $T$ increases. In Fig. S5(c)(d) we show the solutions with UPC $\Delta_{\alpha}(0,0)=0.5\varepsilon_{\Lambda}$ (which was not shown in Fig. S3(c)). In (d), we find the UPC holds well except for the regime $E_{g}/\varepsilon_{\Lambda}>1.5$, $T/\varepsilon_{\Lambda}<0.2$, as the model approaches the atomic limit.
After obtaining $\Delta_{\alpha}(T)$, we use UPC as an approximation and insert $\bar{\Delta}(T)$ into Eq. (II) to compute $D_{s}(T)$. From this, one can further calculate the BKT transition temperature $T_{c}$, as shown in Fig. 4(b) and Fig. 5(b) of MT. Fig. S6 shows the temperature-dependence of $D_{s}$ using the example of the two-band chiral model. The $D_{s}(T)$ curves are very flat at $T<0.3T_{MF}$ and have a finite slope near $T_{MF}$, which is a general feature of flat bands with $s$-wave uniform order parameter. $T_{c}$ is extracted from the $D_{s}(T)$ curves by their intersections with the universal line of $(8/\pi)T$. When $D_{s}(0)/\Delta(0)$ is small, the intersections lie in the flat part of the curves and give $T_{c}\approx(\pi/8)D_{s}(0)$, exhibiting a Uemura relation [57].
VIII S8: $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plots and $T_{c}$ vs $E_{g}$ plots for large $J$ Values
As an example, we discuss these results for the representative value $J=6$. The two cases of weak and strong interaction limits are shown in Fig. S7 and S8, respectively. In both Fig. S7(a) and S8(a), $D_{s}/\Delta$ scales with $J$, except for the non-singular model (the dashed curve in Fig. S8(a)). For the flattened three-band non-singular case, the choice $m_{0}$ that makes the uniform pairing best satisfied depends on $J$; for $J=6$, $m_{0}=0.032\varepsilon_{\Lambda}$, which makes $D_{s}/\Delta$ of this curve not scale with $J$ directly.
Figure S7: The $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling behavior, $T_{c}$ and the averaged uniform pairing $\bar{\Delta}$ vs $E_{g}$ for the two-band model (Eq. 1 of MT), with $J=6$. The conventions are the same as for Fig. 4 of the MT. In (b), $U_{A,B}$ are fixed at $0.78\varepsilon_{\Lambda}$ (blue) and $6.28\varepsilon_{\Lambda}$ (red) for the two weak couplings, respectively. (c) is the zero-temperature averaged uniform pairing $\bar{\Delta}$ vs $E_{g}$. The magnitude of $\bar{\Delta}$ increases a lot due to the large density of states at the edge of the dispersive band of $J=6$.
The large $J$ value makes $D_{s}/\Delta>1$ in most regimes of Fig. S7(a) and S8(a), therefore, the $T_{c}$ enhancement due to the gap closing is mainly contributed by the superconducting gap $\Delta$. For the weak interaction case, one can see the $T_{c}$ enhancement by gap closing for $J=6$ (Fig. S7(b)) is stronger than for $J=1$ (Fig. 4(b) in the MT). The reason is that the edge of the dispersive band of $J=6$ is much flatter than $J=1$, leading to a large increase of the density of states, which further enhances $\Delta$. In order to see for large $J$ the $T_{c}$ is mainly enhanced by $\Delta$ rather than $D_{s}$, the best way is to plot the zero-temperature $\Delta$ vs $E_{g}$, as shown in Fig. S7(c), and compare it with the $T_{c}$ curves in Fig. S7(b). We indeed find they resemble each other. An alternative way is by noticing that $T_{c}/\Delta_{0}$ in Fig. S7(b) is close to $0.5$ as the gap completely closes, since the saturation value is $T_{MF}/\Delta_{0}=0.5$ for flat bands (in the limit $D_{s}\gg\Delta$, the BKT transition temperature $T_{c}=T_{MF}$).
Similarly, in the strong interaction case, we see the $T_{c}$ enhancement of the singular model is no stronger than the non-singular model (Fig. S8(b)), since the enhancement of $\Delta$ by gap closing in both cases are the same (Fig. S8(c)). The three curves in Fig. S7(b) are slightly different because their distinct $D_{s}/\Delta$ behaviors (Fig. S8(a)). The $E_{g}/\Delta_{0}>1$ regime of the curves in Fig. S7(b) also resemble the $\Delta$ vs $E_{g}$ curves in Fig. S8(c), confirming that the $T_{c}$ enhancement is by $\Delta$ rather than $D_{s}$.
Figure S8: The $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling behavior, $T_{c}$ and the averaged uniform pairing $\bar{\Delta}$ vs $E_{g}$ for the flattened three-band model (Eq. 3 of MT, case (ii)), with $J=6$. The conventions are the same as for Fig. 5 of the MT, except here we fix $m=m_{0}=0.032\varepsilon_{\Lambda}$. In (b), $U_{A,B,C}=2.51\varepsilon_{\Lambda}$ (red) and $12.56\varepsilon_{\Lambda}$ (light blue), while the non-singular model curve (dashed purple) is independent of the interactions. In (c), $\bar{\Delta}$ of the three cases ($\Delta_{0}=0.1\varepsilon_{\Lambda}$, $\Delta_{0}=0.5\varepsilon_{\Lambda}$ and non-singular) coincide.